# Investigating a new approach of robustly self-testing two-qubit entangled states

### v2

@@ I Introduction
The most extreme case in device-independent quantum certification is identifying the underlying states and measurements themselves. The so-called self-testing protocol was firstly introduced by Mayer and Yao [12, 13]. Since then, many research results on self-testing have been proposed, ranging from self-testing entangled state [14, 15, 16, 17], incompatible measurements [15, 18, 19], steerable assemblages [20], preparation and measurements [21, 22, 23], entangled measurements [24], and so on (see Ref. [25] for a comprehensive review of self-testing).
@@ III Self-testing of entangled states: Definition and Formulation
$\operatorname{tr}\Big{[}(VE_{a|x}V^{\dagger})\otimes(WE_{b|y}W^{\dagger})\cdot(V\otimes W){\varrho^{\text{AB}}}(V^{\dagger}\otimes W^{\dagger})\Big{]}.$
@@
$\operatorname{tr}$ $\Big{[}(VE_{a|x}V^{\dagger})\otimes(WE_{b|y}W^{\dagger})\cdot(V\otimes W){\varrho^{\text{AB}}}(V^{\dagger}\otimes W^{\dagger})\Big{]}$ (6)
@@ IV Self-testing of the two-qubit maximally entangled state
$\operatorname{tr}_{\rm AB}\Big{\{}U_{\rm\tiny{SWAP}}\Big{[}|0\rangle^{\rm A^{\prime}}\langle 0|\otimes\rho^{\rm AB}\otimes|0\rangle^{\rm A^{\prime}}\langle 0|\Big{]}U_{\rm\tiny{SWAP}}^{\dagger}\Big{\}}$ (17)
@@
$1\displaystyle\frac{A_{2}^{\rm opt}-A_{1}^{\rm opt}A_{2}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 0|\right.+\frac{\openone-A_{1}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 1|\right.$
@@
$1\displaystyle\frac{B_{2}^{\rm opt}-B_{1}^{\rm opt}B_{2}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 0|\right.+\frac{\openone-B_{1}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 1|\right.$
@@ V Self-testing of the two-qubit partially entangled states
$1\displaystyle\frac{A_{2}^{\rm opt}-A_{1}^{\rm opt}A_{2}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 0|\right.+\frac{\openone-A_{1}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 1|\right.$
@@
$1\displaystyle\frac{B_{4}^{\rm opt}-B_{3}^{\rm opt}B_{4}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 0|\right.+\frac{\openone-B_{3}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 1|\right.$

### v3

# Investigating an approach of robustly self-testing two-qubit entangled states

@@ I Introduction
The most extreme case in device-independent (DI) quantum certification is identifying the underlying states and measurements themselves. The so-called self-testing protocol was firstly introduced by Mayers and Yao [12, 13]. Since then, many research results on self-testing have been proposed, ranging from self-testing entangled state [14, 15, 16, 17], incompatible measurements [15, 18, 19], steerable assemblages [20], preparation and measurements [21, 22, 23], entangled measurements [24], and so on (see Ref. [25] for a comprehensive review of self-testing).
@@ III Self-testing of entangled states: Definition and Formulation
$\operatorname{tr}\Big{[}(VE_{a|x}V^{\dagger})\otimes(WE_{b|y}W^{\dagger})(V\otimes W){\varrho^{\text{AB}}}(V^{\dagger}\otimes W^{\dagger})\Big{]}.$
@@
$\operatorname{tr}$ $\Big{[}(VE_{a|x}V^{\dagger})\otimes(WE_{b|y}W^{\dagger})(V\otimes W){\varrho^{\text{AB}}}(V^{\dagger}\otimes W^{\dagger})\Big{]}$ (6)
@@ IV Self-testing of the two-qubit maximally entangled state
$\operatorname{tr}_{\rm AB}\Big{\{}U_{\rm\tiny{SWAP}}\Big{[}|0\rangle^{\rm A^{\prime}}\langle 0|\otimes\rho^{\rm AB}\otimes|0\rangle^{\rm B^{\prime}}\langle 0|\Big{]}U_{\rm\tiny{SWAP}}^{\dagger}\Big{\}}$ (17)
@@
+$1\displaystyle+\frac{A_{2}^{\rm opt}-A_{1}^{\rm opt}A_{2}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 0|\right.+\frac{\openone-A_{1}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 1|\right.$
@@
+$1\displaystyle+\frac{B_{2}^{\rm opt}-B_{1}^{\rm opt}B_{2}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 0|\right.+\frac{\openone-B_{1}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 1|\right.$
@@ V Self-testing of the two-qubit partially entangled states
+$1\displaystyle+\frac{A_{2}^{\rm opt}-A_{1}^{\rm opt}A_{2}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 0|\right.+\frac{\openone-A_{1}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 1|\right.$
@@
+$1\displaystyle+\frac{B_{4}^{\rm opt}-B_{3}^{\rm opt}B_{4}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 0|\right.+\frac{\openone-B_{3}^{\rm opt}}{2}\otimes\left.|1\right\rangle\left\langle 1|\right.$